# The Cayley transform on representations

### v1

@@ Abstract.
The classical Cayley transform is a birational map between a quadratic matrix group and its Lie algebra, which was first discovered by Cayley in 1846. Because of its essential role in both pure and applied mathematics, the classical Cayley transform has been generalized from various perspectives. This paper is concerned with a representation theoretic generalization of the classical Cayley transform. The idea underlying this work is that the applicability of the classical Cayley transform heavily depends on how the Lie group is represented. The goal is to characterize irreducible representations of a Lie group, to which the classical Cayley transform is applicable. To this end, we first establish criteria of the applicability for a general Lie group. If the group is semisimple, we further obtain a geometric condition on the weight diagram of such representations. Lastly, we provide a complete classification for classical complex simple Lie groups and their compact real forms. Except for the previously known examples, spin representations of $\operatorname{Spin}(8)$ are the only ones on our list.
@@ 1. Introduction
    Diagonal matrix group [26, Example 1.20]: $C$ is applicable to $T_{n}$ where $T_{n}\subseteq\operatorname{GL}_{n}(\mathbb{F})$ is the group consisting of all invertible diagonal matrices.
@@
[26, Problem 1.7] Which linear algebraic group can have a generalized Cayley transform?
It is worth remarking that an equivariant birational isomorphism between $G$ and its Lie algebra is not necessarily as useful as the classical Cayley transform $C$ defined in (1). For example, if $G=R_{n}$ is the group of $n\times n$ unipotent matrices whose Lie algebra is $\mathfrak{r}_{n}$ consisting of nilpotent matrices, then
$\exp:\mathfrak{r}_{n}\to R_{n},\quad\exp(u)=\sum_{j=0}^{n-1}\frac{u^{j}}{j!}$
is an equivariant birational isomorphism with inverse $\log:R_{n}\dasharrow\mathfrak{r}_{n}$ defined by $\log(X)=-\sum_{j=1}^{n-1}(I_{n}-X)^{j}/j$. Obviously, computing $\exp(u)$ is much more expensive than evaluating $C(u)$.
@@
Here by Definition 4.1, $C$ is applicable to $\rho$ if there is an open neighbourhood $V$ of 00 in $\mathfrak{g}$ such that $C\circ d\rho(V)\subseteq\rho(G)$. We notice that if $G$ is a linear algebraic group and $\rho:G\to\operatorname{GL}(\mathbb{V})$ is a faithful representation to which $C$ is applicable, then clearly $G$ satisfies the requirement in the LPR problem. This paper is devoted to answering Problem 1.2. First, we establish in Theorem 4.9 the equivalence between the applicability of $C$ to $\rho$ and the power span property of $d\rho(\mathfrak{g})$ (cf. Definition 4.5). Next, we restrict our discussion to semisimple Lie groups in Section 5 and obtain the equivalence between the applicability of $C$ to $\rho$ and the power span property of $d\rho(\mathfrak{h})$ in Theorem 5.2. Here $\mathfrak{h}$ denotes the Cartan subalgebra of $\mathfrak{g}$. Moreover, we prove in Theorem 5.6 that if $\rho$ is irreducible and $C$ is applicable to $\rho$, then the weight diagram of $\rho$ must be the orbit of the highest weight under the action of Weyl group, possibly union with the origin. Lastly, Section 6 is concerned with a complete classification of irreducible representations of classical complex simple Lie groups and their compact real forms, to which $C$ is applicable. Our classification results (cf. Theorems 6.13 and 6.15) indicate that the only irreducible representations to which $C$ is applicable are:
@@
Our results may be recognized as representation theoretic analogues of [26, Theorem 1.31].
@@ 2.1. Real and complex semisimple Lie groups
Let $\mathfrak{g}$ be a real semisimple Lie algebra. We denote its complexification by $\mathfrak{g}^{\mathbb{C}}\coloneqq\mathfrak{g}\otimes\mathbb{C}$. A complexification of a real semisimple Lie group $G$ is the simply-connected complex semisimple Lie group $G^{\mathbb{C}}$ whose Lie algebra is isomorphic to $\mathfrak{g}^{\mathbb{C}}$. A real form of a complex semisimple Lie algebra $\mathfrak{u}$ is a real Lie subalgebra $\mathfrak{u}^{\prime}$ of $\mathfrak{u}$ such that ${\mathfrak{u}^{\prime}}^{\mathbb{C}}\simeq\mathfrak{u}$. Let $U$ be a simply-connected complex semisimple Lie group. A real form of $U$ is a real Lie subgroup $U^{\prime}$ of $U$ such that $U$ is a complexification of $U^{\prime}$.
Given a complex semisimple Lie group $U$ and a representation $\rho:U\to\operatorname{GL}(\mathbb{V})$, the restriction of $\rho$ to a real form $U^{\prime}$ of $U$ is
$\operatorname{Res}(\rho)\coloneqq\rho|_{U^{\prime}}:U^{\prime}\to\operatorname{GL}(\mathbb{V}),\quad\operatorname{Res}(\rho)(g)=\rho(g).$
Similarly, the restriction of a representation $\pi:\mathfrak{u}\to\mathfrak{gl}(\mathbb{V})$ of a complex semisimple Lie algebra $\mathfrak{u}$ to its real form $\mathfrak{u}_{0}$ is defined by $\operatorname{Res}(\pi)(x)=\pi(x)$ for $x\in\mathfrak{u}_{0}$.
Suppose that $\pi:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$ is a representation of a real semisimple Lie algebra $\mathfrak{g}$. The induction of $\pi$ is the representation of $\mathfrak{g}^{\mathbb{C}}$ defined as
$\operatorname{Ind}(\pi):\mathfrak{g}^{\mathbb{C}}\to\mathfrak{gl}(\mathbb{V}),\quad\operatorname{Ind}(\pi)(x_{1}+ix_{2})=\rho(x_{1})+i\rho(x_{2}),\;x_{1},x_{2}\in\mathfrak{g}.$
If $\rho:G\to\operatorname{GL}(\mathbb{V})$ is a representation of a real semisimple Lie group $G$ and $G^{\mathbb{C}}$ is its complexification. The induction of $\rho$ is the representation $\operatorname{Ind}(\rho):G^{\mathbb{C}}\to\operatorname{GL}(\mathbb{V})$ uniquely determined by $\operatorname{Ind}(d\rho):\mathfrak{g}^{\mathbb{C}}\to\mathfrak{gl}(\mathbb{V})$. To be more explicit, we have
$\operatorname{Ind}(\rho):G^{\mathbb{C}}\to\operatorname{GL}(\mathbb{V}),\quad\operatorname{Ind}(\rho)(\exp(x))=\exp(\operatorname{Ind}(d\rho)(x)),\;x\in\mathfrak{g}^{\mathbb{C}}$
where $\exp$ denotes the exponential map of Lie groups.
@@
[13, Theorem 14.18] Let $\mathfrak{g}$ be a semisimple Lie algebra. Suppose that $\Lambda$ (resp. $\Phi$, $\mathcal{W}$, $C$) is the weight lattice (resp. set of roots, Weyl group, fundamental Weyl chamber) of $\mathfrak{g}$. Then we have the following:
@@
    Let $\rho:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$ be the irreducible finite dimensional representation with highest weight $\omega$ and let $\mathbb{V}=\bigoplus_{\lambda\in\mathbb{V}}\mathbb{V}_{\lambda}$ be the weight space decomposition. Then $\{\lambda\in\Lambda:\mathbb{V}_{\lambda}\neq 0\}=\operatorname{Conv}(\mathcal{W}\omega)\cap\left(\omega+\Lambda\right)$, where $\mathcal{W}\omega$ is the $\mathcal{W}$-orbit of $\omega$ and $\operatorname{Conv}(\mathcal{W}\omega)$ denotes the convex hull of $\mathcal{W}\omega$.
@@ 2.2. Symmetric algebraic groups
A linear algebraic group $G\subseteq\operatorname{GL}_{n}(\mathbb{C})$ is called symmetric if $G^{\ast}=G$, where $G^{\ast}\coloneqq\left\{\overline{g}^{\scriptscriptstyle\mathsf{T}}:g\in G\right\}$
@@
[37, Lemma 3.5, Theorem 3.13] Let $G\subseteq\operatorname{GL}_{n}(\mathbb{C})$ be a linear algebraic group. Then $G$ is linearly reductive if and only if there is some $P\in\operatorname{GL}_{n}(\mathbb{C})$ such that $PGP^{-1}$ is a symmetric group.
@@
[33, Section 4.1.2] For each semisimple Lie subalgebra $\mathfrak{g}\subseteq\mathfrak{gl}_{n}(\mathbb{F})$, there exists an irreducible algebraic subgroup $G\subseteq\operatorname{GL}_{n}(\mathbb{F})$ such that $\mathfrak{g}$ is the Lie algebra of $G$. Moreover, if $\mathbb{F}=\mathbb{C}$, then $G$ can be chosen to be connected.
Theorem 2.4 (Polar decomposition).
[37, Theorem 2.12] Let $G\subseteq\operatorname{GL}_{n}(\mathbb{C})$ be a symmetric algebraic group and let $K\coloneqq G\cap U_{n}$ where $U_{n}$ is the group of $n\times n$ unitary matrices. We denote by $\mathfrak{k}$ the Lie algebra of $K$. Then the map $p:K\times i\mathfrak{k}\to G$ defined by $p(k,Z)\coloneqq k\exp(Z)$ is a diffeomorphism. In particular, $K$ is connected if and only if $G$ is an irreducible algebraic variety.
We remark that the equivalence between the connectedness of $K$ and $G$ follows immediately from the diffeomorphism $K\times i\mathfrak{k}\simeq G$. By [37, Corollary 2.14] , the connectedness of $G$ is further equivalent to the irreducibility of $G$ as an algebraic variety.
@@
[37, Corollary 2.20] Let $G$ be a connected symmetric subgroup of $\operatorname{GL}_{n}(\mathbb{C})$ and let $K$ = $G\cap U_{n}$ be as above. Then we have $\mathfrak{g}=\mathfrak{k}\bigoplus_{\mathbb{R}}i\mathfrak{k}$. Moreover, If $T$ is a maximal compact torus of $K$ with Lie algebra $\mathfrak{t}$, then $\mathfrak{t}$ is a Cartan algebra of $K$ and $\mathfrak{k}=\mathrm{Ad}(K)\mathfrak{t}$.
@@ 3. Some basic facts
Property (a) follows immediately from the definition. Suppose that $\pi$ is irreducible and $0\neq\mathbb{V}^{\prime}\subseteq\mathbb{V}$ is a subspace such that $\operatorname{Ind}(\pi)(\mathfrak{g})(\mathbb{V}’)\subseteq\mathbb{V}_{1}$. Then we have
$\pi(\mathfrak{g})(\mathbb{V}^{\prime})=\operatorname{Res}(\operatorname{Ind}(\pi))(\mathfrak{g}))(\mathbb{V}’)=\operatorname{Ind}(\pi)(\mathfrak{g})(\mathbb{V}’)\subseteq\mathbb{V}^{\prime},$
@@
Let $\mathfrak{g}$ be a complex semisimple Lie algebra and let $\mathrm{ad}:\mathfrak{g}\to\mathfrak{gl}(\mathfrak{g})$ be its adjoint representation. There exists an irreducible symmetric algebraic group $G\subseteq\operatorname{GL}(\mathfrak{g})$ whose Lie algebra is equal to $\mathrm{ad}(\mathfrak{g})\simeq\mathfrak{g}$.
@@
The semisimplicity of $\mathfrak{g}$ ensures that $\mathrm{ad}$ is faithful. Hence we have $\mathfrak{g}\simeq\mathrm{ad}(\mathfrak{g})\subseteq\mathfrak{gl}(\mathfrak{g})$. According to Proposition 2.3, there is an irreducible algebraic subgroup $G^{\prime}\subseteq\operatorname{GL}(\mathfrak{g})$ whose Lie algebra is equal to $\mathrm{ad}(\mathfrak{g})$. We claim that $G^{\prime}$ is linearly reductive. Then Lemma 2.2 implies the existence of $P\in\operatorname{GL}(\mathfrak{g})$ such that $G\coloneqq PG^{\prime}P^{-1}$ is the desired group. Therefore, it is left to prove the claim. Let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a finite dimensional rational representation of $G$. Since the Lie algebra of $G$ is isomorphic to $\mathfrak{g}$, $\rho$ induces a representation $\pi:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$. The semisimplicity of $\mathfrak{g}$ implies $\pi=\prod_{j=1}^{m}\pi_{j}:\mathfrak{g}\to\bigoplus_{j=1}^{m}\mathfrak{gl}(\mathbb{V}_{j})$ where $\pi_{j}:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V}_{j})$ is an irreducible representation for each $1\leq j\leq m$. We recall that $\pi$ is induced by $\rho$, thus $\rho$ admits a decomposition $\rho=\prod_{j=1}^{m}\rho_{j}:G\to\prod_{j=1}^{m}\operatorname{GL}(\mathbb{V}_{j})$ where $\rho_{j}:G\to\operatorname{GL}(\mathbb{V}_{j})$ is a representation inducing $\pi_{j}$. Moreover, by construction $\rho_{j}$ is irreducible for each $1\leq j\leq m$. ∎
@@
We notice that $\ker\left(\widetilde{\mathrm{Ad}}\right)=Z\left(\widetilde{G}\right)$, $\ker(p)\subseteq Z\left(\widetilde{G}\right)$ and $p\left(Z\left(\widetilde{G}\right)\right)\subseteq Z(G)=\ker(q)$. We may conclude that $\ker(q\circ p)=p^{-1}(Z(G))=Z\left(\widetilde{G}\right)$. Indeed, the inclusion $Z\left(\widetilde{G}\right)\subseteq p^{-1}(Z(G))$ is obtained by
$Z\left(\widetilde{G}\right)\subseteq p^{-1}\left(p\left(Z\left(\widetilde{G}\right)\right)\right)\subseteq p^{-1}\left(Z(G)\right).$
For the reverse inclusion, we observe that $xyx^{-1}y^{-1}\in\ker(p)$ for each $x\in p^{-1}(Z(G))$ and $y\in\widetilde{G}$. The discreteness of $\ker(p)=\pi_{1}(G)$ and the connectedness of $\widetilde{G}$ imply that $xyx^{-1}y^{-1}=h_{x}\in G$ only depends on $x$. Hence, we have $xyx^{-1}y^{-1}=1$ and $p^{-1}(Z(G))\subseteq Z\left(\widetilde{G}\right)$. The commutativity of the diagram follows immediately. ∎
@@
Let $G$ be a compact semisimple Lie group with Lie algebra $\mathfrak{g}$ and let $\mathfrak{h}$ be the Cartan subalgebra of $\mathfrak{g}$. Then we have $\mathfrak{g}=\mathrm{Ad}(G)(\mathfrak{h})$.
@@
Without loss of generality, we assume that $G$ is connected. By Lemma 3.2, there is a symmetric irreducible algebraic subgroup $G^{\prime}\subseteq\operatorname{GL}\left(\mathfrak{g}^{\mathbb{C}}\right)$ whose Lie algebra is isomorphic to $\mathfrak{g}^{\mathbb{C}}$. Let $U\subseteq\operatorname{GL}\left(\mathfrak{g}^{\mathbb{C}}\right)$ be the subgroup consisting of unitary elements and let $K\coloneqq G^{\prime}\cap U$. By Theorem 2.4, $K$ is connected and $\mathfrak{k}^{\mathbb{C}}\simeq\mathfrak{g}^{\mathbb{C}}$. The uniqueness of the compact form of a semisimple Lie algebra implies $\mathfrak{k}\simeq\mathfrak{g}$. Moreover, by Theorem 2.5 we have $\mathfrak{k}=\mathrm{Ad}(K)(\mathfrak{h}_{K})$ where $\mathfrak{h}_{K}$ is the Cartan subalgebra of $\mathfrak{k}$. Under the isomorphism $\mathfrak{k}\simeq\mathfrak{g}$, Lemma 3.3 implies $\mathfrak{g}=\mathrm{Ad}(G)\mathfrak{h}$. ∎

### v2

@@ Abstract.
The classical Cayley transform is a birational map between a quadratic matrix group and its Lie algebra, which was first discovered by Cayley in 1846. Because of its essential role in both pure and applied mathematics, the classical Cayley transform has been generalized from various perspectives. This paper is concerned with a representation theoretic generalization of the classical Cayley transform. The idea underlying this work is that the applicability of the classical Cayley transform heavily depends on how the Lie group is represented. The goal is to characterize irreducible representations of a Lie group, to which the classical Cayley transform is applicable. To this end, we first establish criteria of the applicability for a general Lie group. If the group is semisimple, we further obtain a geometric condition on the weight diagram of such representations. Lastly, we provide a complete classification for complex simple Lie groups and their compact real forms. Except for the previously known examples, spin representations of $\operatorname{Spin}(8)$ are the only ones on our list.
@@ 1. Introduction
    Diagonal matrix group [31, Example 1.20]: $C$ is applicable to $T_{n}$ where $T_{n}\subseteq\operatorname{GL}_{n}(\mathbb{F})$ is the group consisting of all invertible diagonal matrices.
@@
The above mentioned applications of the Cayley transform is a consequence of the fact that it establishes a correspondence between a Lie group and its Lie algebra. It is noticeable that the Cayley transform can also be defined on the complex plane or some space of operators, which makes it a fundamental ingredient in complex analysis [28, 42, 44] and operator theory [43, 30, 24, 10].
@@
[31, Problem 1.7] Which linear algebraic group can have a generalized Cayley transform?
@@
Here by Definition 4.1, $C$ is applicable to $\rho$ if there is an open neighbourhood $V$ of 00 in $\mathfrak{g}$ such that $C\circ d\rho(V)\subseteq\rho(G)$. We notice that if $G$ is a linear algebraic group and $\rho:G\to\operatorname{GL}(\mathbb{V})$ is a faithful representation to which $C$ is applicable, then clearly $G$ satisfies the requirement in the LPR problem. This paper is devoted to answering Problem 1.2. First, we establish in Theorem 4.9 the equivalence between the applicability of $C$ to $\rho$ and the power span property of $d\rho(\mathfrak{g})$ (cf. Definition 4.5). Next, we restrict our discussion to semisimple Lie groups in Section 5 and obtain the equivalence between the applicability of $C$ to $\rho$ and the power span property of $d\rho(\mathfrak{h})$ in Theorem 5.2. Here $\mathfrak{h}$ denotes the Cartan subalgebra of $\mathfrak{g}$. Moreover, we prove in Theorem 5.6 that if $\rho$ is irreducible and $C$ is applicable to $\rho$, then the weight diagram of $\rho$ must be the orbit of the highest weight under the action of Weyl group, possibly union with the origin. Lastly, Section 6 is concerned with a complete classification of irreducible representations of complex simple Lie groups and their compact real forms, to which $C$ is applicable. Our classification results (cf. Theorems 6.13 and 6.15) indicate that the only irreducible representations to which $C$ is applicable are:
@@
Our results may be recognized as representation theoretic analogues of [31, Theorem 1.31].
Caveats
In this paper, we consider Lie groups and Lie algebras over $\mathbb{R}$ and $\mathbb{C}$. Unless otherwise stated, representations to be discussed are all finite-dimensional complex vector spaces (except for the adjoint representations of Lie groups and Lie algebras), even for real Lie groups and real Lie algebras. A morphism between two real (resp. complex) Lie groups is a smooth (resp. complex analytic) group homomorphism.
@@ 2.1. Real and complex semisimple Lie groups
2.1.1. Complexification
Let $\mathfrak{g}$ be a real semisimple Lie algebra. We denote its complexification by $\mathfrak{g}^{\mathbb{C}}\coloneqq\mathfrak{g}\otimes\mathbb{C}$. A complexification of a real Lie group $G$ [9, Chapitre 3] is a complex Lie group $G^{\mathbb{C}}$ together with a real Lie group morphism $\tau:G\to G^{\mathbb{C}}$ satisfying the following universal property: For any complex Lie group $H$ and any real Lie group morphism $\psi:G\to H$, there exists a unique complex Lie group morphism $\varphi:G^{\mathbb{C}}\to H$ such that $\varphi\circ\tau=\psi$. The complexification of a connected real Lie group $G$ with Lie algebra $\mathfrak{g}$ always exists. If moreover $G$ is simply connected, then the Lie algebra of $G^{\mathbb{C}}$ is isomorphic to $\mathfrak{g}^{\mathbb{C}}$ [9, Chapitre 3,4,5,6].
2.1.2. Real form
A real form of a complex semisimple Lie algebra $\mathfrak{u}$ [15, Lecture 26] is a real Lie subalgebra $\mathfrak{u}^{\prime}$ of $\mathfrak{u}$ such that ${\mathfrak{u}^{\prime}}^{\mathbb{C}}\simeq\mathfrak{u}$. Let $U$ be a connected complex Lie group. A real form of $U$ is a real Lie subgroup $U^{\prime}$ of $U$ such that $U$ is a complexification of $U^{\prime}$ and $\mathfrak{u}^{\prime\mathbb{C}}\simeq\mathfrak{u}$. Real forms of a complex Lie group are not unique in general. In Section 5, we will discuss two important real forms of complex semisimple Lie groups: the split form and the compact form.
2.1.3. Restricted and induced representation
Given a complex semisimple Lie algebra $\mathfrak{u}$ and a representation $\pi:\mathfrak{u}\to\mathfrak{gl}(\mathbb{V})$, the restricted representation of $\pi$ to its real form $\mathfrak{u}_{0}$ is defined by
$\operatorname{Res}(\pi)\coloneqq\pi|_{\mathfrak{u}_{0}}:\mathfrak{u}_{0}\to\mathfrak{gl}(\mathbb{V}),\quad\operatorname{Res}(\pi)(x)=\pi(x).$
Similarly, given a complex semisimple Lie group $U$ and a representation $\rho:U\to\operatorname{GL}(\mathbb{V})$, the restricted representation of $\rho$ to a real form $U_{0}$ of $U$ is
$\operatorname{Res}(\rho)\coloneqq\rho|_{U_{0}}:U_{0}\to\operatorname{GL}(\mathbb{V}),\quad\operatorname{Res}(\rho)(g)=\rho(g).$
Suppose that $\pi:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$ is a representation of a real semisimple Lie algebra $\mathfrak{g}$. The induced representation of $\pi$ is
$\operatorname{Ind}(\pi):\mathfrak{g}^{\mathbb{C}}\to\mathfrak{gl}(\mathbb{V}),\quad\operatorname{Ind}(\pi)(x_{1}+ix_{2})=\pi(x_{1})+i\pi(x_{2}),$
where $x_{1},x_{2}\in\mathfrak{g}$. Since $\mathbb{V}$ is a complex vectors space, $\operatorname{Ind}(\pi)$ is well-defined. If $\rho:G\to\operatorname{GL}(\mathbb{V})$ is a representation of a real semisimple Lie group $G$ and $(G^{\mathbb{C}},\tau)$ is its complexification. The induced representation of $\rho$ is the representation $\operatorname{Ind}(\rho):G^{\mathbb{C}}\to\operatorname{GL}(\mathbb{V})$ uniquely determined by the universal property of $(G^{\mathbb{C}},\tau)$. If moreover $G$ is a real form of $G^{\mathbb{C}}$, then explicitly we have
$\operatorname{Ind}(\rho):G^{\mathbb{C}}\to\operatorname{GL}(\mathbb{V}),\quad\operatorname{Ind}(\rho)(\exp(x))=\exp(\operatorname{Ind}(d\rho)(x)),$
where $x\in\mathfrak{g}^{\mathbb{C}}$ and $\exp:\mathfrak{g}\to G$ is the exponential map.
2.1.4. Weight lattice and irreducible representation
Let $\mathfrak{g}$ be a semisimple Lie algebra. Suppose that $\Lambda$ (resp. $\Phi$, $\mathcal{W}$, $C$) is the weight lattice (resp. set of roots, Weyl group, fundamental Weyl chamber) of $\mathfrak{g}$.
@@ 2.1.4. Weight lattice and irreducible representation
[15, Theorem 14.18] We have the following:
@@
    Let $\rho:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$ be the irreducible finite dimensional representation with highest weight $\omega$ and let $\mathbb{V}=\bigoplus_{\lambda\in\mathbb{V}}\mathbb{V}_{\lambda}$ be the weight space decomposition. Then $\{\lambda\in\Lambda:\mathbb{V}_{\lambda}\neq 0\}=\operatorname{Conv}(\mathcal{W}\omega)\cap\left(\omega+\mathbb{Z}\Phi\right)$, where $\mathcal{W}\omega$ is the $\mathcal{W}$-orbit of $\omega$ and $\operatorname{Conv}(\mathcal{W}\omega)$ denotes the convex hull of $\mathcal{W}\omega$.
@@ 2.2. Symmetric algebraic groups
A complex linear algebraic group $G\subseteq\operatorname{GL}_{n}(\mathbb{C})$ is called linearly reductive [45, Section 3.1.3] if all the finite-dimensional regular representations are completely reducible . A complex linear algebraic group $G\subseteq\operatorname{GL}_{n}(\mathbb{C})$ is called symmetric [45, Section 3.1.3] if $G^{\ast}=G$, where $G^{\ast}\coloneqq\left\{\overline{g}^{\scriptscriptstyle\mathsf{T}}:g\in G\right\}$. It is worth remarking that linear reductivity is independent of embeddings, whereas the definition of a symmetric group depends on the specific embedding of $G$ into $\operatorname{GL}_{n}(\mathbb{C})$. Lemma 2.2 below reveals that linearly reductive groups and symmetric groups are two sides of the same coin.
@@
[45, Lemma 3.5, Theorem 3.13] Let $G\subseteq\operatorname{GL}_{n}(\mathbb{C})$ be a linear algebraic group. Then $G$ is linearly reductive if and only if there is some $P\in\operatorname{GL}_{n}(\mathbb{C})$ such that $PGP^{-1}$ is a symmetric group.
Theorem 2.3 (Polar decomposition).
[45, Theorem 2.12] Let $G\subseteq\operatorname{GL}_{n}(\mathbb{C})$ be a symmetric algebraic group and let $K\coloneqq G\cap U_{n}$ where $U_{n}$ is the group of $n\times n$ unitary matrices. We denote by $\mathfrak{k}$ the Lie algebra of $K$. Then the map $p:K\times i\mathfrak{k}\to G$ defined by $p(k,Z)\coloneqq k\exp(Z)$ is a diffeomorphism. In particular, $K$ is connected if and only if $G$ is an irreducible algebraic variety.
A subgroup $H$ of an algebraic group $G$ is called an irreducible algebraic subgroup if $H$ is an irreducible subvariety of $G$.
@@
[38, Section 4.1.2] For each semisimple Lie subalgebra $\mathfrak{g}\subseteq\mathfrak{gl}_{n}(\mathbb{F})$, there exists an irreducible algebraic subgroup $G\subseteq\operatorname{GL}_{n}(\mathbb{F})$ such that $\mathfrak{g}$ is the Lie algebra of $G$. Moreover, if $\mathbb{F}=\mathbb{C}$, then $G$ can be chosen to be connected in the Euclidean topology.
We remark that the equivalence between the connectedness of $K$ and $G$ follows immediately from the diffeomorphism $K\times i\mathfrak{k}\simeq G$. By [45, Corollary 2.14] , the connectedness of $G$ is further equivalent to the irreducibility of $G$ as an algebraic variety.
@@
[45, Corollary 2.20] Let $G$ be a connected symmetric subgroup of $\operatorname{GL}_{n}(\mathbb{C})$ and let $K$ = $G\cap U_{n}$ be as above. Then we have $\mathfrak{g}=\mathfrak{k}\bigoplus_{\mathbb{R}}i\mathfrak{k}$. Moreover, If $T$ is a maximal compact torus of $K$ with Lie algebra $\mathfrak{t}$, then $\mathfrak{t}$ is a Cartan algebra of $K$ and $\mathfrak{k}=\mathrm{Ad}(K)\mathfrak{t}$.
@@ 3. Some basic facts
Property (a) follows immediately from the definition. Suppose that $\pi$ is irreducible and $0\neq\mathbb{V}^{\prime}\subseteq\mathbb{V}$ is a subspace such that $\operatorname{Ind}(\pi)(\mathfrak{g})(\mathbb{V}’)\subseteq\mathbb{V}$. Then we have
$\mathbb{V\subseteq}\pi(\mathfrak{g})(\mathbb{V}^{\prime})=\operatorname{Res}(\operatorname{Ind}(\pi))(\mathfrak{g}))(\mathbb{V}’)=\operatorname{Ind}(\pi)(\mathfrak{g})(\mathbb{V}’)\subseteq\mathbb{V}^{\prime},$
@@
Let $\mathfrak{g}$ be a complex semisimple Lie algebra and let $\mathrm{ad}:\mathfrak{g}\to\mathfrak{gl}(\mathfrak{g})$ be its adjoint representation. There exists an irreducible linearly reductive algebraic group $G\subseteq\operatorname{GL}(\mathfrak{g})$ whose Lie algebra is isomorphic to $\mathrm{ad}(\mathfrak{g})\simeq\mathfrak{g}$. By a change of coordinates of $\mathfrak{g}$, one can furthermore require $G$ to be symmetric.
@@
The semisimplicity of $\mathfrak{g}$ ensures that $\mathrm{ad}$ is faithful. Hence we have $\mathfrak{g}\simeq\mathrm{ad}(\mathfrak{g})\subseteq\mathfrak{gl}(\mathfrak{g})$. According to Proposition 2.4, there is an irreducible algebraic subgroup $G^{\prime}\subseteq\operatorname{GL}(\mathfrak{g})$ whose Lie algebra is equal to $\mathrm{ad}(\mathfrak{g})$. We claim that $G^{\prime}$ is linearly reductive. Then Lemma 2.2 implies the existence of $P\in\operatorname{GL}(\mathfrak{g})$ such that $G\coloneqq PG^{\prime}P^{-1}$ is the desired group. Therefore, it is left to prove the claim. Let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a finite dimensional rational representation of $G$. Since the Lie algebra of $G$ is isomorphic to $\mathfrak{g}$, $\rho$ induces a representation $\pi:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$. The semisimplicity of $\mathfrak{g}$ implies $\pi=\prod_{j=1}^{m}\pi_{j}:\mathfrak{g}\to\bigoplus_{j=1}^{m}\mathfrak{gl}(\mathbb{V}_{j})$ where $\pi_{j}:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V}_{j})$ is an irreducible representation for each $1\leq j\leq m$. We recall that $\pi$ is induced by $\rho$, thus $\rho$ admits a decomposition $\rho=\prod_{j=1}^{m}\rho_{j}:G\to\prod_{j=1}^{m}\operatorname{GL}(\mathbb{V}_{j})$ where $\rho_{j}:G\to\operatorname{GL}(\mathbb{V}_{j})$ is a representation inducing $\pi_{j}$. Moreover, by construction $\rho_{j}$ is irreducible for each $1\leq j\leq m$.
By choosing a basis for $\mathfrak{g}$, we may identify $\operatorname{GL}(\mathfrak{g})$ with $\operatorname{GL}_{n}(\mathbb{C})$, where $n=\dim\mathfrak{g}$. If $G\subseteq\operatorname{GL}(\mathfrak{g})$ is linearly reductive, then Lemma 2.2 implies that $PGP^{-1}\subseteq\operatorname{GL}(\mathfrak{g})$ is symmetric for some $P\in\operatorname{GL}(\mathfrak{g})$. As a consequence, the change of coordinates of $\mathfrak{g}$ determined by $P$ renders $G$ a symmetric group. ∎
@@
Notice that $\ker\left(\widetilde{\mathrm{Ad}}\right)=Z\left(\widetilde{G}\right)$, $\ker(p)\subseteq Z\left(\widetilde{G}\right)$ and $p\left(Z\left(\widetilde{G}\right)\right)\subseteq Z(G)=\ker(q)$. We may conclude that $\ker(q\circ p)=p^{-1}(Z(G))=Z\left(\widetilde{G}\right)$. The commutativity of the diagram follows immediately. ∎
@@
Let $G$ be a compact semisimple Lie group with Lie algebra $\mathfrak{g}$. We fix a root system for $G$ and denote by $\mathfrak{h}$ the Cartan subalgebra of $\mathfrak{g}$. Then we have $\mathfrak{g}=\mathrm{Ad}(G)(\mathfrak{h})$.
@@
Without loss of generality, we assume that $G$ is connected. Since any two Cartan subalgebras of $\mathfrak{g}$ are conjugated by the action of $G$, we are free to choose the one stated in Theorem 2.5. By Lemma 3.2, there is a symmetric irreducible algebraic subgroup $G^{\prime}\subseteq\operatorname{GL}\left(\mathfrak{g}^{\mathbb{C}}\right)$ whose Lie algebra is isomorphic to $\mathfrak{g}^{\mathbb{C}}$. Let $U\subseteq\operatorname{GL}\left(\mathfrak{g}^{\mathbb{C}}\right)$ be the subgroup consisting of unitary elements and let $K\coloneqq G^{\prime}\cap U$. By Theorem 2.3, $K$ is connected and $\mathfrak{k}^{\mathbb{C}}\simeq\mathfrak{g}^{\mathbb{C}}$. The uniqueness of the compact form of a semisimple Lie algebra implies $\mathfrak{k}\simeq\mathfrak{g}$. Moreover, by Theorem 2.5 we have $\mathfrak{k}=\mathrm{Ad}(K)(\mathfrak{h}_{K})$ where $\mathfrak{h}_{K}$ is the Cartan subalgebra of $\mathfrak{k}$. Under the isomorphism $\mathfrak{k}\simeq\mathfrak{g}$, Lemma 3.3 implies $\mathfrak{g}=\mathrm{Ad}(G)\mathfrak{h}$. ∎